# Can Transformers Do Enumerative Geometry?

### v1

@@ 6.3.2. Large genus limit and abductive reasoning
This theorem illustrates that, as the genus $g$ grows large, the intersection numbers grow factorially as $(2g)!$, with a specific exponential behavior modulated by the constant $A=2/3$. Recently, another approach for computing the large-genus asymptotics of intersection numbers was proposed by [eynard2023resurgentlargegenusasymptotics], based on the resurgent analysis of the $n$-point functions of such intersection numbers. With this approach, the authors extended Aggarwal’s results by computing all subleading corrections:
@@
The process of identifying different components and fixing parameters (such as $A=2/3$ or the subleading corrections) can be approached using abductive reasoning, framing it as a hypothesis-testing problem. Abduction involves proposing plausible explanations for observed data, which is similar to hypothesizing different theories to determine which best aligns with the AI agent’s understanding of the data, such as the embedding vector space in models like Transformers. In hypothesis testing, one considers two hypotheses: the null hypothesis $H_{0}$ and the alternative hypothesis $H_{1}$. The goal is to decide whether the observed data sufficiently support the null hypothesis. Specifically, given a set of observations, the aim is to maximize the probability of making a discovery while controlling for a prescribed level of type I error.
In this context, one can investigate whether the AI agent has correctly understood parameters, such as the constant $A=2/3$ in the asymptotic formula, and whether this understanding can be recovered as an inverse problem. Thus, the important question we want to answer in computational enumerative geometry is: In scenarios where parameters in asymptotic formulas are unknown, can we use abduction to infer and provide evidence for their potential values? To address this question, we employ correlational and causal methods based on the Transformer’s internal understanding of the data, which can be applied to various enumerative problems. First, we check the reasoning of DynamicFormer by deciphering the model’s understanding of the parameters of the asymptotic form of the $\psi$-class intersection numbers, which are not explicitly provided to the model. The parameters of interest to be inferred and rediscovered are as follows:
  • The exponential growth $A$.
  • The first few subleading asymptotic polynomials $\alpha_{1},\alpha_{2},\ldots$ [guo2022largegenusasymptoticspsiclass, eynard2023resurgentlargegenusasymptotics].
@@
One can claim that there is evidence that the model’s embedding vector space $p_{g,n}(\langle\bm{d}\rangle_{g,n}|\bm{B},\bm{d})$ encodes the parameter $A=2/3$ and the subleading polynomials if there is a correspondence between the embeddings and these parameters. A prominent method that attempts to shed light on this correspondence is probing [conneau2018cramsing], also known as diagnostic prediction [hupkes2020compositionality]. Under this methodology, one trains a simple model (e.g., kernel Ridge Regression [pml1Book]) as a probe to predict some desired information from the latent representations of the Transformer model. High prediction performance is interpreted as evidence that the information is encoded in the vector space representation. The efficacy of the encoding can be gauged by the amount of variance in these parameters that can be explained by $p_{g,n}(\langle\bm{d}\rangle_{g,n}|\bm{B},\bm{d})$, quantified by the $\mathrm{R}^{2}$ value of the probe. To quantify the uncertainty of our abductive reasoning through probing, we perform the conformal prediction procedure [shafer2008tutorial] as a statistical uncertainty quantification approach to obtain uncertainty bands around the predictions.
For $A=2/3$, we set up the experiment with a grid of alternative hypotheses. Specifically, we probe a discrete value space for the numerator of $A$, $\mathrm{num}={1,\ldots,10}$, and the denominator, $\mathrm{denom}={1,\ldots,10}$, which covers 90 candidates (with $A=1$ excluded as a trivial hypothesis). We find evidence that DynamicFormer encodes $A$ as $2/3$ with the highest $\mathrm{R}^{2}$ and lowest uncertainty, as shown in figure 7.
@@
Figure 7. The model’s internal embedding’s predictive power of the exponential growth constant of $\psi$-class intersection numbers with hypothesis testing, gauged by the $\mathrm{R}^{2}$ and conformal uncertainty.
We perform the same analysis for the first few terms of the subleading series $\alpha_{k}$ with $k=0,1,2,3$. We found that with $90\%$ coverage, the Transformer’s vector space representation can predict the coefficients with an $\mathrm{R}^{2}=0.98$ and a conformal prediction interval of ±$\pm 531$. This interval is only about $5.4\%$ of the standard deviation of the coefficients, implying that the prediction interval is relatively narrow compared to the overall variability in the data. This suggests that nearly $98\%$ of the variance in the subleading coefficients is encoded in and predictable from the model’s internal vector space.
If we perform Principal Component Analysis (PCA) on the hidden embedding of each sample, $p_{g,n}$, we observe in figure 8 that DynamicFormer shows hierarchical representations of intersection numbers across both the genus $g$ and the number of marked points $n$. This provides evidence that the model has learned a recursive structure within the data. The PCA plots highlight that the two principal components with the highest variance form circular patterns, reflecting the model’s internal representation of amplitudes.
Figure 8. The first two principle components of model’s hidden states across different genera (left) and number of marked points (right).
@@
So far, we have shown evidence of the possibility of recovering the exponential growth constant and the first few subleading terms from the AI agent. Now, we want to analyze the internal mechanisms and the inner causal understanding of the model via causal tracing [vig2020]. Transformers can be viewed as a causal graph, where information is propagated from the input to the output through a network of hidden states. Originally, causal tracing was introduced to quantify the storage and transfer of facts within the intermediate variables of Transformers [meng2022locating, feng2024language, wang2024grokked], which is not our focus here. We seek to use causal tracing to understand the model’s decision-making process, specifically identifying which input variables and modalities are more influential when predicting the intersection of $\psi$-classes. In other words, we want to uncover how the AI agent comprehends the underlying mathematics. We will refer to an input modality as $m[i]=\set{n[i],\bm{B}[i],\bm{d}[i]}$ for each input instance $i$. There are then two steps:
  1. (1)
@@
    Clean run. This run records the model’s prediction on a regular input for a fixed genus $g$. With this run, we record auto-correlation of the predicted intersection numbers across different number of marked points.
  2. (2)
@@
    Counter-factual intervention. In a counter-factual intervention [baron2023explainable], we input modified instances for each modality into the model to observe how these changes affect the predictions. These counter-factuals involve incorrectly assigning features to instances. For example, this could mean incorrectly associating the partition $\bm{d}=(29,10,5,3,1,0)$ with $n=3$ instead of $n=6$. Previous research on factual associations in Transformers has explored methods such as adding noise to inputs [meng2022locating], which creates redundant distribution shifts [zhang2024best], and exchanging tokens [feng2024language]. Our approach is similar to the latter method, focusing on how changes in input tokens affect model predictions. Specifically, for the modality of interest (while keeping the other modalities unchanged), we replace the input instances with random alternatives of the same genus. This replacement results in a misassignment of features to a sample, which alters the target prediction, i.e., $m[i]\rightarrow m[j]$ where $i\neq j$, to obtain $p_{g,n}(\langle\bm{d}\rangle_{g,n}|\mathrm{do}(m[i]))$. The purpose of this intervention is to study how the model relates features to instances and to understand the underlying mathematical relationships between them. This analysis allows us to attribute cause and effect between interventions and the amplitudes. The counter-factual analysis provides causal evidence on the role of each input modality. The difference observed between the clean and intervened runs, as measured by $\mathrm{R}^{2}$, the autocorrelation of the intersection numbers, and the $\mathrm{R}^{2}_{\textup{probe}}$ of the probe for the exponential growth constant $A=2/3$, characterizes the causal impact of the modality of interest on the $\psi$-class intersection numbers.
Table 5. Causal Strength for Different Modalities
$\mathrm{do}(m[i])$ $\mathrm{R}^{2}$ $\mathrm{R}^{2}_{\text{probe}}$
No Intervention $0.98$ $0.93$
$n$ $-12.6$ $0.53$
$\bm{B}$ $0.54$ $-3.62$
$\bm{d}$ $-52.2$ $0.44$
@@
Figure 9. Autocorrelation plot across different number of marked points $n$ for clean (leftmost) and intervened runs for $n$, $\bm{B}$, and $\bm{d}$.
@@
What causal reasoning is the model learning? We compute the $\mathrm{R}^{2}$ and $\mathrm{R}^{2}_{\textup{probe}}$ over the samples with $g=15$ to measure causal strength. As a result, the $\mathrm{R}^{2}$ of the intersection of $\psi$-classes is heavily affected when using counter-factual instances for the number of marked points $n$ and partitions $\bm{d}$. This indicates a strong causal connection between the partitions and the intersection numbers. This is also evident from the autocorrelation plot figure 9, where the correlation between the intersection numbers is completely lost after disrupting the binding between instances and their partitions. The presence of such a strong causal relationship for $n$ and $\bm{d}$ is not surprising, but their mild impact on $\mathrm{R}^{2}_{\textup{probe}}$ is an interesting discovery. Despite a weak causal connection between $\bm{B}$ and the predicted intersection numbers, there is evidence that the exponential growth constant, $A=2/3$, is causally connected to the quantum Airy structure initial datum $\bm{B}$. This suggests that the model’s internal reasoning (hidden states) of the $\bm{B}$ initial datum encodes this parameter. Meanwhile, the weak causal impact of $\bm{B}$ on the intersection numbers results in only a mild effect on their autocorrelation, which was unexpected.
7. Conclusion and Discussions
@@
While the capabilities of Large Language Models and Transformers, as their primary building blocks, have advanced significantly to encompass complex tasks across various domains, it remains an open question whether they can be effectively utilized for abstract mathematical reasoning and knowledge discovery. In this paper, we address a fundamental question: Can Transformers perform and learn enumerative geometry and topological recursion? To answer this question, we investigated a deep but theoretically well-studied enumerative problem in algebraic geometry, specifically the computation of $\psi$-class intersection numbers. Given that these amplitudes exhibit recursive behavior, capturing this behavior presents a non-trivial challenge for Transformers. We analyze the ability of Transformers to tackle such a task in a zero-shot setting, where the AI model is trained only on genus $g\leq 14$ and is asked to predict intersection numbers for $g_{\textup{ID}}=14$ and $g_{\textup{OOD}}=[15,16]$. As a result, we introduced DynamicFormer, a multi-modal Transformer-based model. DynamicFormer takes the quantum Airy structure datum $\bm{B}$, the partition $\bm{d}$, and the global genus $g$ and number of marked points $n$ as input modalities to predict the $\psi$-class intersection numbers. Throughout our experiments, we applied various techniques and modifications to Transformers to enhance their performance. We also identified the crucial role of non-linear activation functions in recursive tasks and introduced a new activation function, Dynamic Range Activator (DRA), which improves prediction precision in both ID and OOD settings.
In the second part of the results, we posed a straightforward question: How exactly does the AI agent perform enumerative geometry? To answer this question, we conducted a series of correlational, conformal, and causal interoperability analyses.
Firstly, by examining the internal vector representation of the model, we discovered that the Transformer is indeed learning Virasoro constraints. Specifically, we found strong evidence that the model has learned at least the Dilaton equations. This is particularly fascinating because no such constraints or conditions were imposed on the model; nonetheless, it managed to pick up and (re)discover them. This finding is intriguing because it suggests that in cases where such relations and equations are not known but are suspected to exist, a pre-trained Transformer could potentially aid in a human-machine collaboration by providing hints and evidence for these relations in a purely data-driven manner.
@@
Afterwards, we further explored whether it is possible to perform some form of abductive reasoning and hypothesis testing to estimate the parameters of asymptotic formulas for intersection numbers, or more generally, for quantum Airy structure amplitudes. Often, there is an expectation for such asymptotic growths, which then leads to a process of conjecture building and deductive reasoning to derive the final formula. In this work, we advocate that using the Transformer’s internal representation and understanding of the data, one can provide informed guesses and even reject alternative hypotheses about candidate parameters. To achieve this, we conducted a grid search on a parameter of the asymptotic formula for $\psi$-class intersection numbers, specifically the exponential growth constant, based on the Transformer’s hidden representation. Using probing techniques and conformal prediction, we demonstrated that this information is indeed encoded in the model’s internal representation and provided strong evidence for recovering its value. We also performed the same analysis (without any alternative hypotheses) on the first few terms of the subleading terms. We found that nearly $98\%$ of the variance in the subleading terms is encoded in and can be predicted with high confidence from the model’s internal vector space.
@@
In the end, we analyzed the internal decision-making procedure of DynamicFormer using causal inference. By applying the causal tracing method with counterfactual interventions, we aimed to shed light on how various input modalities—namely, the partitions, quantum Airy structure, and number of marked points—are causally responsible for the model’s understanding of the intersection numbers of $\psi$-classes and the parameters for large genus.
7.1. Limitations and Future Directions
It is important to address some key limitations of this work and studies in this area. One of the central messages of this research is that simply predicting or classifying a mathematical object with machine learning is not sufficient. Even with a highly accurate model, uncertainties in predictions always exist, and it is crucial to account for them. A prediction without proper uncertainty estimation is not reliable, especially in mathematics, where observations and computations must be both robust and precise.
In our work, we aimed to enhance reliability and explainability, but uncertainties in predictions and knowledge discovery still persist. For instance, if the underlying quantum Airy structure changes, the model may struggle to generalize. One potential solution would be to incorporate symbolic embeddings [dascoli2022deep, meidani2024snip] of the quantum Airy structures, modulated by numerical values, to develop a foundational model for intersection numbers. This approach could facilitate the study of relationships between different theories and ultimately enable the symbolic prediction of closed-form formulas for asymptotic behaviors of intersection numbers.
Another limitation of our work is the performance of uncertainty estimation, where we employed conformal prediction. While introducing this type of uncertainty estimation to the community is a step forward, our current coverage is clearly insufficient, indicating that our uncertainty bands are not yet fully reliable. Future efforts should explore more robust and appropriate conformal methods [xu2024conformal] for uncertainty quantification.
A fascinating avenue for future research is to explore recent developments in intersection numbers [eynard2023newformula]. This new approach proposes a formula that involves sums over partitions of combinatorial factors and reveals that certain partitions with specific conditions have vanishing coefficients in their decomposition into elementary symmetric polynomials. Remarkably, empirical observations suggest that even more partitions have vanishing coefficients than the trivial ones. This unexpected result points to a deeper, hidden structure in intersection numbers, where the internal representation of the Transformer model could be very beneficial in examining these new results and uncovering the underlying patterns.

### v3

@@ 5.1.1 Large Genus Asymptotic and Abductive Reasoning
This theorem, proven by Aggarwal (2021), illustrates that, as the genus $g$ grows large, the intersection numbers grow factorially as $(2g)!$ with a specific exponential behavior modulated by the constant $A=2/3$. Recently, another approach for computing the large genus asymptotics of intersection numbers was proposed by Eynard et al. (2023). The strategy of this proof is based on a resurgent analysis of the $n$-point functions of $\psi$-class intersection numbers, which are computed via determinantal formulae. The determinantal formula is a consequence of the integrability property of the intersection numbers, namely KdV. With this approach, the authors extended Aggarwal’s results by computing all subleading corrections:
@@
Abductive Reasoning: Identifying components and parameters such as $A=2/3$ or the subleading corrections in Equation 5.4 can also be approached through abductive reasoning, framing the task as hypothesis testing. Abduction involves proposing plausible explanations for observed data, to determine which best aligns with the network’s understanding of the data. In this context, we investigate how the network can infer the parameters like the constant $A=2/3$ in the asymptotic formula, treating it as an inverse problem. Thus, an important question we want to answer is: In scenarios where parameters in asymptotic formulas are unknown, can we use abduction to infer and provide evidence for their potential values? To address this, we first examine the reasoning of DynamicFormer by deciphering its understanding of the asymptotic parameters of the intersection numbers, which are not explicitly provided to the model. The parameters we aim to infer and rediscover are the exponential growth $A$ and the first few subleading asymptotic polynomials $\alpha_{1},\alpha_{2},\ldots$ (Guo & Yang, 2022; Eynard et al., 2023).
@@
One can claim that there is an evidence that the model’s embedding vector space $\bm{x}_{g,n}(\langle\bm{d}\rangle_{g,n}|\bm{B},\bm{d})$ encodes the parameter $A=2/3$ and the subleading polynomials if there is a correspondence between the embeddings and these parameters. A prominent method that attempts to shed light on this correspondence is probing (Conneau et al., 2018), also known as diagnostic prediction (Hupkes et al., 2020). Under this methodology, one trains a linear or non-linear model as a probe to predict any desired information from the latent representations of the network. Formally, we aim to evaluate how well the network’s hidden representations encode the fundamental parameter $A$ by predicting the target function $I_{j}$ associated with each input $j$. Let $\bm{x}^{j}_{g,n}\in\mathbb{R}^{k}$ be the hidden state for input $j$, and $I_{j}$ be defined based on approximate conjectural estimation of Equation 5.4. To map the hidden representations $\bm{x}^{j}_{g,n}$ to the target values $I_{j}$, we employ both linear and non-linear (MLP) probes $f$. The probe $f$ is trained, using a conformal prediction procedure (Shafer & Vovk, 2008) to quantify the statistical uncertainty, by minimizing the squared error between the predicted and actual values of $I_{j}$, $f=\arg\min_{\theta}\sum_{j}(f_{\theta}(\bm{x}^{j}_{g,n})-I_{j})^{2}.$ High prediction performance is interpreted as an evidence that the information is encoded in the Transformer’s enumerative world model. The efficacy of the encoding is gauged by the coefficient of determination ($\mathrm{R}^{2}$) of the probe.
For $A=2/3$, we set up the experiment with a grid of alternative hypotheses. Specifically, we probe a discrete value space for the numerator of $A$, $\mathrm{num}={1,\ldots,10}$, and the denominator, $\mathrm{denom}={1,\ldots,10}$, which covers 90 candidates (with $A=1$ excluded as a trivial hypothesis). It is expected from the Wentzel–Kramers–Brillouin (WKB) method applied to the underlying Airy quantum curve that $A$ would be a period on the associated Riemann surface. For this specific enumerative problem, periods are in $\mathbb{Q}$. We use both linear and non-linear probes to see if the network’s internal representation for $A=2/3$ has a linear form or not. The higher and more precise performance of the non-linear probe to the linear one suggests that $A$ could be encoded non-linearly, as shown in Figure 3. We perform the same analysis for the first few terms of the subleading series $\alpha_{k}$ with $k=0,1,2,3$. We found that with $90\%$ coverage, the Transformer’s vector space representation can predict the coefficients with an $\mathrm{R}^{2}=0.98$. Notably, the linear probe yields a lower value $\mathrm{R}^{2}=0.63$. This suggests that the network’s internal representation of the polynomiality phenomenon (Guo & Yang, 2022; Eynard et al., 2023) does not have a simple linear form, but a non-linear representation.
@@
Figure 3: The network’s internal representation’s linear (left) and non-linear (right) predictive power of the exponential growth constant of $\psi$-class intersection numbers.
@@
Causal Tracing: Originally, causal tracing (Vig et al., 2020) was introduced to quantify information storage and transfer within Transformer components (Meng et al., 2022; Feng & Steinhardt, 2024; Wang et al., 2024), which is not our focus here. Instead, we use causal tracing to analyze the model’s decision-making, specifically identifying which input modalities influence the prediction of $\psi$-class intersections. In other words, we aim to uncover how the network causally comprehends the underlying mathematics. We define an input modality as $m[i]=\set{n[i],\bm{B}[i],\bm{d}[i]}$ for each instance $i$. There are then two steps:
  1. 1.
@@
    Clean run. This run records the model’s prediction on a regular input. With this run, we record auto-correlation of the predicted intersection numbers across different number of marked points.
  2. 2.
@@
    Counter-factual intervention. We perform counterfactual interventions (Baron, 2023) by modifying instances in one modality while keeping others unchanged to observe how these changes affect the model’s predictions. This involves incorrectly assigning features to instances—for example, associating the partition $\bm{d}=(29,10,5,3,1,0)$ with $n=3$ instead of $n=6$. Previous research on factual associations in Transformers has explored methods such as adding noise to inputs (Meng et al., 2022), which creates redundant distribution shifts (Zhang & Nanda, 2024), and exchanging tokens (Feng & Steinhardt, 2024). Our approach is similar to the latter method, focusing on how changes in input tokens affect model predictions. In our case, for the modality of interest (while keeping the other modalities intact), we replace the input instances with random alternatives of the same genus. This replacement results in a miss-assignment of features to a sample, which alters the target prediction, i.e., $m[i]\rightarrow m[j]$ where $i\neq j$, to obtain $\bm{x}_{g,n}(\langle\bm{d}\rangle_{g,n}|\mathrm{do}(m[i]))$. This allows us to study how the model relates features to instances and to attribute cause and effect between interventions and the intersections. The difference observed between the clean and intervened runs is measured by $\mathrm{R}^{2}$, the autocorrelation of the intersection numbers, and the $\mathrm{R}^{2}_{\textup{probe}}$ of the linear probe for the exponential growth constant $A=2/3$.
Table 3: Causal Strength for Different Modalities
No Intervention $n$ $\bm{B}$ $\bm{d}$
$\mathrm{R}^{2}$ $0.96$ $-12.6$ $0.54$ $-52.2$
$\mathrm{R}^{2}_{\text{probe}}$ $0.83$ $0.52$ $-2.77$ $0.43$
@@
Figure 4: Autocorrelation plot across different number of marked points $n$ for clean (leftmost) and intervened runs (from left to right) for $n$, $\bm{B}$, and $\bm{d}$ respectively.
@@
What causal reasoning is the model learning? $\mathrm{R}^{2}$ of the intersection of $\psi$-classes is heavily affected when using counter-factual instances for the number of marked points $n$ and partitions $\bm{d}$. This indicates a strong causal connection between the partitions and the intersection numbers. This is also evident from the autocorrelation plot Figure 4, where the correlation between the intersection numbers is completely lost after disrupting the binding between instances and their partitions. The presence of such a strong causal relationship for $n$ and $\bm{d}$ is not surprising, but their mild impact on $\mathrm{R}^{2}_{\textup{probe}}$ is an interesting discovery. Despite a weak causal connection between $\bm{B}$ and the predicted intersection numbers, there is an evidence that the exponential growth constant, $A=2/3$, is causally connected to the quantum Airy structure initial datum $\bm{B}$. This suggests that the network’s hidden states of the $\bm{B}$ initial datum encodes this parameter. Meanwhile, the weak causal impact of $\bm{B}$ on the intersection numbers results in only a mild effect on their autocorrelation. This is expected since $\bm{B}$ contributes linearly, shown in Equation 2.4, while the dependence in $n$ and $d$ are factorial.
6 Conclusion
@@
In this work, we tried to address a fundamental question: Can Transformers perform and learn enumerative geometry and topological recursion? To answer, we investigated a deep enumerative problem in algebraic geometry, specifically the computation of $\psi$-class intersection numbers. As a result, we introduced DynamicFormer, a multi-modal Transformer-based model, to predict the $\psi$-class intersection numbers. We analyzed the ability of the network to tackle such a task in a zero-shot setting. Capturing the recursive behavior of these invariants poses a non-trivial challenge for Transformers. Throughout our experiments, we identified the crucial role of non-linear activation functions and introduced a new one, Dynamic Range Activator (DRA), which improves prediction precision of recursive maps. One important message of this work is that merely predicting or classifying a mathematical object with ML is insufficient. A prediction without proper uncertainty estimation is unreliable, especially in mathematics. We enhance reliability by incorporating conformal uncertainty estimation and explainability for knowledge discovery.
@@
In the second part of the results, we asked a simple question: How exactly does the network perform enumerative geometry? To answer this, we conducted a series of correlational, conformal, and causal interoperability analyzes. Firstly, by examining the internal vector space of the model, we discovered that the network is learning Virasoro constraints without them being imposed on the model. This finding is intriguing because it suggests that in cases where such relations and equations are unknown but are suspected to exist, a pre-trained Transformer could potentially aid in a human-machine collaboration by providing hints and evidence for these relations in a purely data-driven manner. We further explored whether it is possible to perform some form of abductive reasoning and hypothesis testing to estimate the parameters of asymptotic form for intersection numbers. Often, there is an expectation for such asymptotic growths that leads to a process of conjecture building and deductive reasoning to derive the final formula. In this work, we advocate that by using the network’s internal understanding of the data, one can provide informed guesses and even reject alternative hypotheses about candidate parameters. To achieve this, we conducted a grid search on the exponential growth constant of the asymptotic limit for $\psi$-class intersection numbers, based on the network’s hidden representation. Using probing techniques, we showed that this information is encoded non-linearly in the model’s internal representation. We also performed the same analysis on the first few terms of the subleading polynomial terms. We found that these subleading terms are also encoded in a non-linear manner in the model’s hidden space.
@@
In the end, we analyzed the internal decision-making procedure of DynamicFormer using causal inference. By applying the causal tracing method with counterfactual interventions, we aimed to shed light on how various input modalities—namely, the partitions, quantum Airy structure data, and number of marked points—are causally responsible for the model’s understanding of $\psi$-class intersection numbers and their large genus parameters.